OpenStudy (anonymous):

Indicate in standard form the equation of the line through the given points. K(6, 4), L(-6, 4)

OpenStudy (anonymous):

to find the slop would i have to do the (y2-y1)/ (x2-x1) which would look like (4-4)/ (-6-6)= 0/12

OpenStudy (anonymous):

y-y1=m(x-x1) 4-4=m(6-(6) 0=12m m=0 (You have a horizontal line).

OpenStudy (anonymous):

*6-(-6)